# Balanced cross-Kerr coupling for superconducting qubit readout

### v1

@@ Abstract
where $\mathrm{erfc}$ is the complementary error function and $T_{1}$ is the transmon lifetime. The coherent state approximated assignment errors are computed from performing Monte-Carlo simulations using QuTiP’s mcsolve solver [61, 62] to get the resonator’s field $\alpha_{e/g}(t)$. The Hamiltonian and the parameters used for the simulations are the same as what was used for the heterodyne readout simulations. The collapse operator is $\sqrt{\kappa}\hat{a}$ and we use $512$ trajectories for the Monte-Carlo simulations.
@@
As previously mentioned, we employ a two-step pulse, stopping the drive at $t=20$ ns, letting the resonator to decay rapidly thereafter with a rate $\kappa/2\pi=2\chi_{z}/2\pi\simeq 20$ MHz. To quantify the QNDness, we track the population of each transmon level $i_{t}$ by calculating $P_{i_{t}}(t)=\mathrm{Tr}\left[\sum_{n_{r}}|\overline{i_{t},n_{r}}\rangle\langle\overline{i_{t},n_{r}}|\hat{\rho}(t)\right]$. The transmon populations as a function of time for initial states $|\overline{0_{t},0_{r}}\rangle$ and $|\overline{1_{t},0_{r}}\rangle$ are shown in Fig. S3(b) and (c), respectively. From Fig. S3(a) we identify the time at which the resonator population drops below $10^{-2}$ and use this to evaluate the population of the ground and first excited states of the transmon to characterize the quantum nondemolition (QND) nature of the readout. For these parameters, the QND fidelity was found to be $99.89\%$ when starting in the ground state and $99.71\%$, when starting in the first excited state.
While weak resonances in the transmon-resonator spectrum are often disregarded when analyzing measurement-induced transitions, they are crucial for characterizing leakage rates. The extent to which the transmon leaks depends significantly on the degree of hybridization between the transmon and resonator states [19, 21]. To fully quantify this hybridization, in Fig. S4, we plot the overlap between the bare and dressed states of the transmon-resonator system for various resonator photon numbers. We find that for dispersive readout, the overlap decreases rapidly even at moderate photon numbers, particularly when the system starts in the first excited state. In contrast, for junction readout, the overlap exhibits only minimal variation with increasing photon numbers, suggesting that junction readout results in less leakage compared to dispersive readout, as the system’s state deviates less from its initial qubit state during the readout process.
@@
In this section, we outline the quantum optimal control method used to enhance the readout fidelity and speed, as shown in Fig. 4(c). These simulations are based on Ref. [63] which makes uses of the open source library dynamiqs [64]. We also use the evosax library [65] to efficiently explore the large optimization parameter space with the Covariance Matrix Adaptation (CMA) evolution strategy.
@@
As discussed in Ref. [63], the readout pulse is divided into $1$ ns time bins, representing the parameters to be optimized, and filtered through a Gaussian envelope with a $500$ MHz bandwidth. Furthermore, the maximum amplitude is limited to $\epsilon_{\textrm{max}}/2\pi=700$ MHz whilst ensuring that the photon population does not exceed $\bar{n}_{\textrm{max}}=40$ to remain consistent with our previous simulations (see Fig. 4 and Sec. S5).
@@
The cost function consists of a weighted sum of 4 components, as summarized in Table S1: the inverse of the signal-to-noise ratio (SNR), the maximum resonator population, the maximum amplitude and the pulse smoothness. The optimization is performed for integration times between 10 and 30 ns. Although the pulse shapes vary slightly with integration time, all optimized pulses resemble a two-step pulse and reach the maximum resonator population, see Fig. S5 (b) and (c). Crucially, as shown in Fig. S5 (a), the optimized pulses outperform the analytical pulse described in Sec. S5 at all integration times, reaching a readout fidelity of $99.99$% in $t_{m}\simeq 18\>\textrm{ns}$. Even when accounting for a finite $T_{1}$ of $120\>\mu s$ and a reduced measurement efficiency of $\eta=0.5$, the optimal pulse achieves $99.99$% readout fidelity in $t_{m}\simeq 28\>\textrm{ns}$.
@@
Figure S5: (a) Assignment error obtained from the coherent state approximation (light blue, see Fig. 4(c)) and quantum optimal control (dark blue). Solid lines correspond to the ideal efficiency case of $\eta=1$, while dashed lines correspond to reduced efficiency of $\eta=0.5$, based on state-of-the-art values. Red circles indicate when both finite lifetime ($T_{1}=120\>\mu s$) and reduced readout efficiency ($\eta=0.5$) are taken into account. For both methods, the parameters are identical to those in Fig. 4. (b) Optimal pulse shape obtained for an integration time of $t_{m}=20$ ns and (c) the corresponding resonator population.

### v2

@@ Abstract
where $\mathrm{erfc}$ is the complementary error function and $T_{1}$ is the transmon lifetime. The coherent state approximated assignment errors are computed from performing Monte-Carlo simulations using QuTiP’s mcsolve solver [Qutip_1, Qutip_2] to get the resonator’s field $\alpha_{e/g}(t)$. The Hamiltonian and the parameters used for the simulations are the same as what was used for the heterodyne readout simulations. The collapse operator is $\sqrt{\kappa}\hat{a}$ and we use $512$ trajectories for the Monte-Carlo simulations.
@@
Final resonator population $\sum_{i=g,e}\operatorname{ReLU}\left(\operatorname{Tr}\left[\hat{a}^{\dagger}\hat{a}\hat{\rho}_{i}(t_{\rm{total}})\right]\right)$ $10^{2}$ -
@@
As previously mentioned, we employ a two-step pulse, stopping the drive at $t=20$ ns, letting the resonator decay rapidly thereafter with a rate $\kappa/2\pi=2|\chi_{z}|/2\pi\simeq 20$ MHz. To quantify the QNDness, we track the population of each transmon level $i_{t}$ by calculating $P_{i_{t}}(t)=\mathrm{Tr}\left[\sum_{n_{r}}|\overline{i_{t},n_{r}}\rangle\langle\overline{i_{t},n_{r}}|\hat{\rho}(t)\right]$. The transmon populations as a function of time for initial states $|\overline{0_{t},0_{r}}\rangle$ and $|\overline{1_{t},0_{r}}\rangle$ are shown in Fig. S4(b) and (c), respectively. From Fig. S4(a) we identify the time at which the resonator population drops below $10^{-2}$ and use this to evaluate the population of the ground and first excited states of the transmon to characterize the quantum nondemolition (QND) nature of the readout. For these parameters, the QND fidelity was found to be $99.89\%$ when starting in the ground state and $99.71\%$, when starting in the first excited state. Furthermore, the leakage from the computational states to the second excited state was found to be $0.045\%$ and $0.0014\%$ for the ground and excited state, respectively.
While weak resonances in the transmon-resonator spectrum are often disregarded when analyzing measurement-induced transitions, they are crucial for characterizing leakage rates. The extent to which the transmon leaks depends significantly on the degree of hybridization between the transmon and resonator states [MIST_2, Dumas2024]. To fully quantify this hybridization, in Fig. S5, we plot the overlap between the bare and dressed states of the transmon-resonator system for various resonator photon numbers. We find that for dispersive readout, the overlap decreases rapidly even at moderate photon numbers, particularly when the system starts in the first excited state. In contrast, for junction readout, the overlap exhibits only minimal variation with increasing photon numbers, suggesting that junction readout results in less leakage compared to dispersive readout, as the system’s state deviates less from its initial qubit state during the readout process.
@@
[…]
These simulations build on Ref. [gautier2024optimal] and make use of the open-source library dynamiqs [guilmin2025dynamiqs]. To efficiently explore the large optimization parameter space for readout, we employ the evosax library [evosax2022github] together with the Covariance Matrix Adaptation (CMA) evolution strategy. For the reset optimization, we use the optimistix library [optimistix2024] with the BFGS algorithm, ensuring rapid convergence to short reset times.
@@
As discussed in Ref. [gautier2024optimal], the readout pulse is divided into $1$ ns time bins, representing the parameters to be optimized, and filtered through a Gaussian envelope with a $294$ MHz bandwidth. Furthermore, the maximum amplitude is limited to $\epsilon_{\textrm{max}}/2\pi=700$ MHz whilst ensuring that the photon population does not exceed $\bar{n}_{\textrm{max}}=40$ to remain consistent with our previous simulations (see Fig. 4 and Sec. S5).
@@
Focusing first on optimizing the readout fidelity, the cost function consists of weighted sum of 4 components, as summarized in Table S1: the inverse of the signal-to-noise ratio (SNR), the maximum resonator population, the maximum amplitude, and the pulse smoothness. The optimization is performed for integration times between 10 and 30 ns. Although the pulse shapes vary slightly with integration time, all optimized pulses resemble a two-step pulse and reach the maximum resonator population. Crucially, as shown in Fig. S6 (a), the optimized pulses outperform the analytical pulse described in Sec. S5 at all integration times, reaching a readout fidelity of 99.9999.99% in $t_{m}\simeq 18\>\textrm{ns}$. Even when accounting for a finite $T_{1}$ of $120\>\mu s$ and a measurement efficiency of $\eta=0.5$, the optimal pulse achieves 99.9999.99% readout fidelity in $t_{m}\simeq 28\>\textrm{ns}$.
[…]
@@
Figure S6: (a) Assignment error obtained from the coherent state approximation (light blue, see Fig. 4(c)) and quantum optimal control (dark blue). Solid lines correspond to the ideal efficiency case of $\eta=1$, while dashed lines correspond to reduced efficiency of $\eta=0.5$, based on state-of-the-art values. Red circles indicate when both finite lifetime ($T_{1}=120~\mu s$) and reduced readout efficiency ($\eta=0.5$) are taken into account. For both methods, the parameters are identical to those in Fig. 4. (b) Optimal pulse shape obtained for a total readout time of $t_{\rm{total}}=75$ ns including resonator reset, and (c) the corresponding resonator population. Dashed vertical lines indicate $t_{m}=20$ ns, the time at which the pulse is optimized to reset the resonator.
S8 Removing the flux loop
Figure S7: Circuit implementations of junction readout without a flux loop. (a) Floating transmon coupled to a $\lambda/4$ readout resonator. (b) Grounded transmon coupled to a $\lambda/2$ readout resonator. (c) Single-junction design in which the transmon nonlinearity is entirely inherited from the coupling junction. In contrast to configurations (a) and (b), the circuit in (c) does not feature any parasitic low-frequency modes.
The lumped-element circuit of Fig. 1(a) contains a loop through which a flux can be applied. This can be used to tune the qubit but can also lead to dephasing. In this section we discuss three possible designs for implementing junction readout without the flux loop. The circuit designs are shown in Fig. S7. The designs in panels (a) and (b) rely on breaking the flux loop by either floating the transmon or by using a $\lambda/2$ resonator. These approaches result in an extra mode in the circuit whose consequences we discuss below. In contrast, the design of (c) that breaks the flux loop without introducing an additional mode.
S8.1 Floating transmon design
Consider the circuit in Fig. S7(a). The total circuit Hamiltonian can now be written as
$\hat{H}=\hat{H}_{\text{t,r}}+\hat{H}_{\text{cm}}+\hat{H}_{\text{int}},$ (S12)
where $\hat{H}_{\text{t,r}}$ describes the Hamiltonian of the transmon-resonator system, see Eq. (1) in the main text. Moreover, $\hat{H}_{\text{cm}}$ and $\hat{H}_{\text{int}}$ represents the Hamiltonian of the additional mode and its interaction with the rest of the system, respectively. Below, we investigate the consequences of this additional mode. For simplicity we focus on the regime where $C_{1}\ll C_{0}$, such that the transmon and the additional mode reduced flux variables are given by $\hat{\varphi_{t}}=\hat{\varphi}_{1}-\hat{\varphi}_{0}$ and $\hat{\varphi}_{cm}=(C_{1}/C_{0})\hat{\varphi}_{1}+\hat{\varphi}_{0}\simeq\hat{\varphi}_{0}$  [^1] This simplification also ensures a fair comparison with the grounded case since it maintains similar strong coupling between the transmon and the readout resonator. Alternatively, if $C_{0}\simeq C_{1}$, the charge–charge coupling is reduced by a factor of two, as expected from floating the qubit, and the cross-Kerr interaction is similarly reduced by a factor of four relative to the asymmetric case.. The conclusion that follow remain unchanged if the condition $C_{1}\ll C_{0}$ is not satisfied.
The presence of this additional mode—often referred to as the “common mode” (cm)—in a floating transmon is not unique to junction readout. However, if the transmon is capacitively coupled to the resonator, as in the standard dispersive readout, the common mode is of no consequence. Indeed, its Hamiltonian consists only of a kinetic charging energy term, i.e., $\hat{H}_{\text{cm}}=4E_{C_{cm}}(\hat{n}_{cm}-n_{g_{cm}})^{2}$, and it therefore behaves as a free particle. Moreover, for the same reason, the Hamiltonian of the common mode commutes with the charge operator which is the relevant qubit-resonator coupling operator in the case of dispersive readout where the interaction with the common mode reads
$\hat{H}_{\text{int,cap.}}=J_{cm,t}\hat{n}_{cm}\hat{n}_{t}+J_{cm,r}\hat{n}_{cm}\hat{n}_{r}.$ (S13)
This interaction is also of no consequence for the transmon and the resonator, i.e., it acts as a renormalization of the effective gate charge acting on the transmon. On the other hand, because of the galvanic coupling between the transmon and the readout resonator in junction readout, the common mode acquires a potential-like term, i.e., $\hat{H}_{\text{cm}}=4E_{C_{cm}}(\hat{n}_{cm}-n_{g_{cm}})^{2}-E_{J_{c}}\cos\hat{\varphi}_{cm}$, and non-trivial coupling to both the transmon and readout resonator modes,
$\begin{split}&\hat{H}_{\text{int,jr}}=J_{cm,t}\hat{n}_{cm}\hat{n}_{t}+J_{cm,r}\hat{n}_{cm}\hat{n}_{r}\\ &-E_{J_{c}}\left[\cos(\hat{\varphi}_{cm}+\hat{\varphi}_{t}-\hat{\varphi}_{r})-\cos(\hat{\varphi}_{cm})-\cos(\hat{\varphi}_{t}-\hat{\varphi}_{r})\right].\end{split}$ (S14)
The last two cosine terms in the second line are present to avoid double-counting contributions already included in $\hat{H}_{\mathrm{t,r}}$ and $\hat{H}_{\text{cm}}$. This coupling can result in new multiphoton resonances, something which can lower the ionization critical photon number for the transmon mode.
Fig. S8(c) shows the critical photon number as a function of the resonator frequency $\omega_{r}$ and the ratio $C_{s}/C_{0}$ obtained for the coupled transmon, common mode, and resonator system using branch analysis. In this analysis, the common mode is assumed to initially be in its ground state. Here, $C_{s}$ is the transmon shunt capacitance and $C_{0}$ is the capacitance to ground, see Fig. S7(a), both of which can be chosen at fabrication time. We keep the shunt capacitor $C_{s}$ of the transmon fixed such that the transmon charging energy remains around $E_{C_{t}}/2\pi\simeq 300$ MHz, similar to the grounded transmon case studied in the main text. All other parameters are identical to those in Sec. S2, with common-mode transmon coupling $J_{cm,t}\simeq-JC_{r}/C_{0}$ which depends on the ground capacitance $C_{0}$. Fig. S8(a) compares the critical photon number in the floating design with $C_{s}/C_{0}=0.1$ to the grounded case, and Fig. S8(b) shows the corresponding ratio. We observe that in most instances floating the transmon lowers the critical photon number compared to the grounded case. This is similar to spectator-induced state transitions recently studied in strongly driven multimode systems [singh:2025, connolly2025, dai2025, Hoyau:2024, benhayounekhadraoui2025].
Interestingly, from Fig. S8(c) we find that decreasing the ratio $C_{s}/C_{0}$ increases the critical photon number. This is likely due to the common-mode frequency $\omega_{cm}$ being pushed downward, along with reductions in the coupling strengths $J_{cm,r}\simeq J(C_{s}/C_{0})$ and $J_{cm,t}$. However, reducing $C_{s}/C_{0}$ even more can push the common mode into a frequency range where it becomes thermally populated, which can lead to additional dephasing of the transmon mode [Bertet:2005, Gambetta:2006, Clerk:2007, Sears:2012].
Figure S8: (a) Critical photon number, and (b) the ratio between the critical photon numbers for floating and grounded transmon designs, for various resonator frequencies $\omega_{r}$. Here the ratio $C_{s}/C_{0}=0.1$. (c) Critical photon number of the floating transmon case for varying $C_{s}/C_{0}$ ratios and resonator frequencies $\omega_{r}$. The contours of several common mode frequencies are shown in dashed yellow lines. The coupling strength $J_{cm,t}$ between the common mode and the transmon at the purple, orange and green stars are $J_{cm.t}/2\pi\simeq-23.36,\>-41.32$, and 59.29-59.29 MHz, respectively. The parameters used here are the same as in Sec. S2.
Importantly, we emphasize that the challenges introduced by floating the qubits are not inherent to the junction readout circuit itself but rather a general consequence of using any coupling other than capacitive coupling. Indeed, with inductively coupled readout schemes–or inductive or junction-based two-qubit couplers–this extra mode acquires a potential-like term, leading to similar considerations, see for instance Refs. [Ma2024, chakraborty2025].
S8.2 Using a $\lambda/2$ readout resonator
Figure S9: (a) Circuit implementation where the coupling junction is located at an arbitrary position $x$ along a $l=\lambda/2$ readout resonator. This setup can be modeled as the junction connected to two transmission line segments of length $x$ and lxl-x, respectively as shown in (b). (c) Equivalent circuit-model in which the junction is coupled to an effective impedance given by the parallel combination of the input impedances of the left and right transmission line segments, denoted $Z_{L}(\omega)$ and $Z_{R}(\omega)$.
An alternative way to eliminate the flux loop is to use a $\lambda/2$ readout resonator, as illustrated in Fig. S9(a). The circuit again maps to a floating configuration, but now it is the LC resonator rather than the transmon that is floated, see Fig. S7(b). To show this, consider the input impedance of the transmission-line (TL) resonator as seen by the coupling junction. For the moment, we consider the case where the coupling junction is located at the far end of TL resonator, i.e., at distance $l=\lambda/2$ from the open end. Below, we analyze the general case of coupling at arbitrary position inside the TL.
The input impedance seen by the coupling junction is given by [Pozar2011MicrowaveEngineering]
$Z_{\mathrm{in}}(\omega)=-iZ_{0}\tan^{-1}(\omega l/v),$ (S15)
where $Z_{0}$ and $v$ are the characteristic impedance and phase velocity of the TL, respectively. The relevance of $Z_{\mathrm{in}}(\omega)$ is that its poles identify the resonances of the readout circuitry, while the associated residues set each mode’s impedance—and thus its coupling strength to the qubit—and at the same time accounting for the distributed nature of the circuit [Nigg2012, Solgun2014, Solgun2015, Solgun2019, Labarca2024]. Using the fraction expansion
$\tan^{-1}(z)=\frac{1}{z}+\sum_{n=1}^{+\infty}\frac{2z}{z^{2}-n^{2}\pi^{2}},\>\>z\in\mathbb{C}\setminus\{n\pi,\>n\in\mathbb{Z}\},$ (S16)
yields the Foster decomposition
$Z_{\mathrm{in}}(\omega)=\frac{1}{i\omega C_{0}}+\sum_{n=1}^{\infty}\frac{i\omega C_{r}^{-1}}{\omega_{n}^{2}-\omega^{2}},$ (S17)
where $\omega_{n}=n\pi v/l=n\omega_{r}$, and $C_{r}=\pi/(2Z_{0}\omega_{r})$. The fundamental mode of a $\lambda/2$ resonator thus has an impedance $Z_{r}=2Z_{0}/\pi<Z_{0}$. Crucially, Eq. S17 shows that a $\lambda/2$ TL is equivalent to a collection of harmonic oscillators (second term) in series with a ground capacitance $C_{0}=2C_{r}$ (first term). Fig. 1(b) corresponds to keeping only the fundamental resonance $\omega_{r}$, i.e., the term $n=1$ in the sum. This approximation holds as long as the operating frequency obeys $\omega\ll 2\omega_{r}$.
Importantly, the ground capacitance $C_{0}$ leads to the same considerations discussed in Sec. S8.1: it introduces a low-frequency degree of freedom that couples to both the transmon and the readout mode. Indeed, defining $\hat{\varphi}_{r}=\hat{\varphi}_{1}-\hat{\varphi}_{0}$, and $\hat{\varphi}_{cm}=-\hat{\varphi}_{0}$, the circuit Hamiltonian of Fig. S7(b) takes the same form as Eq. S12, with circuit parameters given by $E_{C_{cm}}=e^{2}/(2C_{0})=E_{C_{r}}/2$, $J_{cm,t}\simeq J/2$, $J_{cm,r}\simeq-J(C_{s}/2C_{r})$. Consequently, the critical photon number for measurement-induced state transitions displays the same features as in Fig. S8.
We now consider the more general case where the coupling junction is located at an arbitrary position $x$ along the $l=\lambda/2$ resonator as shown in Fig. S9(a) and (b). Below, we show that the circuit can still be mapped to a floating configuration, as in Fig. S7 (b), but with the key difference that the mode impedance $Z_{r}$ now depends on the coupling position $x$.
In this configuration, the coupling junction is effectively connected to two open-ended transmission-line segments of lengths $x$ and lxl-x, respectively, as illustrated in Fig. S9(b). Denoting their input impedances as $Z_{L}(\omega)$ and $Z_{R}(\omega)$, the coupling junction sees a total impedance given by the parallel combination of the two, that is,
$Z_{\mathrm{in}}(\omega)=\frac{Z_{L}(\omega)Z_{R}(\omega)}{Z_{L}(\omega)+Z_{R}(\omega)},$ (S18)
see Fig. S9(c). Here, $Z_{L}(\omega)=-iZ_{0}\tan^{-1}(\omega(l-x)/v)$ and $Z_{R}(\omega)=-iZ_{0}\tan^{-1}{(\omega x/v)}$ with $0<x<l$. We assume a homogeneous TL, meaning both segments have the same characteristic impedance $Z_{0}$ and phase velocity $v$.
As before, we expand $Z_{\mathrm{in}}$ in its canonical Foster form,
$Z_{\mathrm{in}}(\omega)=\frac{A_{0}}{i\omega}+\sum_{n=1}^{\infty}\frac{i\omega A_{n}}{\omega_{n}^{2}-\omega^{2}},$ (S19)
to extract the relevant quantities. The residue of the zero-frequency pole,
$A_{0}(x)=\lim_{\omega\to 0}[i\omega Z_{\mathrm{in}}(\omega)],$ (S20)
determines the effective ground capacitance $C_{0}$, while the impedance of the readout mode is set by the residue of the first finite pole,
$A_{1}(x)=\lim_{\omega\to\omega_{1}}[2i(\omega-\omega_{1})Z_{\mathrm{in}}(\omega)].$ (S21)
Using the analytical expressions for $Z_{L}(\omega)$ and $Z_{R}(\omega)$, we find that $A_{0}(x)=Z_{0}v/l=C_{0}^{-1}$, which is independent of the coupling position and matches the value obtained when the coupling is at the end of the resonator. Moreover, from Eq. S18, the poles of $Z_{\mathrm{in}}(\omega)$ fall into three categories: (i) poles of $Z_{L}$ located at $\pi v/x, n+n\in\mathbb{N}_{+}$ (ii) poles of $Z_{R}$ at $n\pi v/(l-x)$, with $n\in\mathbb{N}_{+}$ and (iii) frequencies satisfying the current conservation $Z_{L}(\omega)=-Z_{R}(\omega)$ which occur at $\pi v/l=n\omega_{r}, n+n\in\mathbb{N}_{+}$. Thus, the lowest pole corresponds to $\omega_{r}$, as expected. A straightforward calculation yields the position-dependent impedance of the readout mode,
$Z_{r}(x)=\frac{2Z_{0}}{\pi}\frac{1}{\cos^{2}(\frac{\pi(l-x)}{l})},$ (S22)
with $Z_{r}(x)=A_{1}(x)/\omega_{r}$. This expression shows that the impedance is minimized at the end of the resonator, where $Z_{r}(x=l)=2Z_{0}/\pi<Z_{0}$, and diverges at the center $x=l/2$, as expected.
For completeness, we briefly comment on the case of a $\lambda/4$ readout resonator, where one end is now shorted. In this configuration, there is no zero-frequency pole, i.e., $A_{0}=0$, reflecting the absence of a ground capacitance. Physically, this corresponds to the lack of a floating island and thus the absence of a low-frequency mode when the resonator is coupled galvanically to the qubit. Of course, this comes at the cost of introducing a flux loop in the presence of both $E_{J}$ and $E_{J_{c}}$, see the next section for a single-junction design. Moreover, the position-dependent impedance of the readout mode is $Z_{r}(x)=(4Z_{0}/\pi)\cos^{2}[\pi(l-x)/(2l)]$. Unlike the $\lambda/2$ case, the mode impedance is maximal when the coupling junction is placed at the end $x=l$, taking the value $Z_{r}(x=l)=4Z_{0}/\pi>Z_{0}$. Importantly, $Z_{r}$ can be reduced by moving the coupling point away from the end.
S8.3 Single junction design
Figure S10: (a) Cross-Kerr coupling strength, (b) resonator self-Kerr, and (c) qubit frequency for various resonator impedances $Z_{r}$ and coupling junction size $E_{Jc}$, for the single junction design. The resonator frequency was set to $\omega_{r}/2\pi=11$ GHz. Grayed out region corresponds to where the dispersive approximation is invalid.
The circuit of Fig. S7(c) removes the transmon junction. In this case, the transmon’s Josephson energy is set entirely by the coupling junction $E_{J_{c}}$. Crucially, in this configuration there are no flux loop and no extra modes, and the circuit is reminiscent of the arm qubit proposed in Ref. [kline:2025]. Fig. S10 shows the cross-Kerr, resonator self-Kerr, and qubit frequencies as functions of the coupling junction energy $E_{J_{c}}$ and the resonator impedance $Z_{r}$. As with the circuit of Fig. 1(a), large cross-Kerr with minimal self-Kerr is readily possible. Although $E_{J_{c}}$ now sets both the coupling strength and the qubit frequency, thereby potentially reducing parameter flexibility, we find this is not limiting, particularly for $\omega_{q}/2\pi\lesssim 5~\text{GHz}$.
We now explore two regimes of this design: one where $E_{Jc}$ is chosen such that the qubit frequency is in the typical range of $\omega_{q}/2\pi=4\sim 6$ GHz, and a second where the qubit frequency is below $\omega_{q}/2\pi=2$ GHz, while the impedance was set to $Z_{r}=18\>\Omega$ in both cases.
Figure S11: Gate-charge averaged critical photon number for the single-junction design, shown for (a) $\omega_{q}/2\pi\simeq 4.63$ GHz and (b) $\omega_{q}/2\pi\simeq 1.65$ GHz, as a function of the readout resonator frequency. The resonator impedance was set to $Z_{r}=18\>\Omega$ for both cases. Vertical lines indicate multiphoton resonances up to a 44-photon process, with darker lines corresponding to fewer-photon processes. Note the difference in the x-axis ranges between panels (a) and (b).
In the first case, we choose $E_{Jc}/2\pi=11$ GHz, with $E_{C}/2\pi=275$ MHz, corresponding to a qubit frequency $\omega_{q}/2\pi=4.63$ GHz. Figure S11(a) shows the critical photon number versus resonator frequency obtained for these parameters. Vertical lines mark the locations of multiphoton (up to $4$ photon) resonances that couple the transmon’s ground and first-excited states to higher levels. We find that the critical photon number is large for several ranges of resonator frequencies. With suitable optimization, this single-junction design performs comparably to a standard transmon over the typical qubit-frequency range.
We now focus on the alternative case where the parameters are chosen such that $\omega_{q}/2\pi\lesssim 2$ GHz. For our simulations, we choose $E_{Jc}/2\pi=4$ GHz and $E_{C}/2\pi=100$ MHz, which ensures a large $E_{J}/E_{C}$ ratio whilst maintaining an anharmonicity of $\alpha/2\pi\sim 111$ MHz. Due to the lower qubit frequency, all higher transitions (e.g., $\omega_{03}$ and $\omega_{14}$) also lie at relatively low frequencies. Consequently, numerous resonances occur in their vicinity, leading to low critical photon numbers, see Fig. S11(b). However, beyond a resonator frequency of approximately $\omega_{r}/2\pi=7$ GHz, the critical photon number increases sharply. This behavior mirrors the high-frequency readout of transmons recently reported in Refs. [connolly2025, kurilovich:2025], where the resonator frequency was placed an order of magnitude or more above the qubit frequency. Such a large detuning reduces the likelihood of accidental frequency collisions and thereby suppresses multiphoton transitions. Crucially, Refs. [kurilovich:2025, connolly2025] rely on a charge–charge coupling based dispersive readout, so the large qubit–resonator detuning necessarily comes at the cost of a small dispersive shift. In principle, one could compensate the large detuning by increasing the transmon–resonator coupling capacitance. In practice, however, the dispersive shift remains small as a larger capacitive coupling results in a lower anharmonicity which is not favorable.
These limitations can be mitigated with the single-junction design of Fig. S7(c) since there is no trade off between the cross-Kerr coupling strength and detuning. In addition, the anharmonicity is no longer restricted by a large capacitive coupling, though it may still be constrained by the $E_{J_{c}}/E_{C}$ ratio.
In summary, the single junction design offers similar performance to the standard junction readout design without the inherent flux loop or additional modes in the system. Furthermore, we have shown that high frequency readout of transmons with junction readout may be attractive for high QND readout without the limitations found in the case of dispersive readout.